OpenStudy (firejay5):

17. Construction: Carlos is planning to build a grain bin with a radius of 15 ft. He reads that the recommended slant of the roof is 25 degrees. He wants the roof to overhang the edge of the bin by 1 ft. What should the length x be? Give your answer in feet and inches. Show work and explain answer to get medal for patronage! :D

OpenStudy (firejay5):

@satellite73

OpenStudy (anonymous):

just make a triangle out of it from the center of the circle to the outside.  you know the length of the base leg of the right triangle and the angle it makes. \[\cos \theta=adjacent/hypotenuse \] \[hypotenuse = \frac{ adjacent }{ \cos \theta } = 16/\cos20\]

OpenStudy (firejay5):

Wait why is it tangent instead of cosine or sine????

OpenStudy (anonymous):

you are looking for the hypotenuse actually so you would want to use cosine.

OpenStudy (firejay5):

@nikato why did you take that off

OpenStudy (nikato):

oh, i did it wrong, its actually cos. you know the adjacent and hypothenuse so CAH